OpenStudy (anonymous):

Determine if the Mean Value Theorem for Integrals applies to the function f(x) = x3 − 9x on the interval [−1, 1]. If so, find the x-coordinates of the point(s) guaranteed to exist by the theorem

OpenStudy (zzr0ck3r):

Is the function continuous on the interval? Is the interval closed?

OpenStudy (anonymous):

um it is continuous and i do believe it is closed

OpenStudy (zzr0ck3r):

Ok, so the theorem applies.

OpenStudy (anonymous):

oh really, okay great that was actually the part i was most unsure of

OpenStudy (zzr0ck3r):

Now we need to look at \[F(b) - F(a)=F'(c)(b-a) \] and we have \(F(x) = \int_a^xf(t) dt\) where \(a=-1\) and \(b=1\).

OpenStudy (anonymous):

okay so then \[ \int\limits_{-1}^{1}x^3 -9x\]

OpenStudy (anonymous):

so is c the x-coordinate im looking for?

OpenStudy (zzr0ck3r):

\(\int_{-1}^1(t^3-9t) dt = 2(c^3-9c)\) Correct.

OpenStudy (anonymous):

okay so im getting 2/3 fro F(b) - F(a)

OpenStudy (anonymous):

do i just solve for c now?

OpenStudy (zzr0ck3r):

hmm cubic around the origin shuold even out on that interval

OpenStudy (zzr0ck3r):

but yes, solve for \(c\). I get \(0=2c(c^2-9)\)

OpenStudy (anonymous):

oh woops i see why i got 2/3 , i put t^2 instead of t^3

OpenStudy (zzr0ck3r):

\(-3\) is not in our interval.

OpenStudy (anonymous):

are you sure i should leave it as zero and not 8.5

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

okay nvm sorry iv been solving a lot of area and volume problems like these in which area and volume cant be negative so i am told to interpret negatives as positives put i guess it doesn't apply here :)

OpenStudy (anonymous):

couldn't c = 0

OpenStudy (anonymous):

\[0 = 2c(c^2-9)\] \[0 = 2c\] \[0 = c\]

OpenStudy (zzr0ck3r):

\(0=2c(c^2-9)\\\dfrac{0}{2c}=c^2-9\\0=c^2-9\\c^2=9\)

OpenStudy (anonymous):

c = 3 , 0

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

so is my answer 0 because 3 is not in the given interval?

OpenStudy (zzr0ck3r):

yes ...sorry

OpenStudy (anonymous):

okay great, thanks for the help :D

OpenStudy (zzr0ck3r):

I think at some point i said \(3\in [-1,1]\) doooh. deleted it...

OpenStudy (zzr0ck3r):

np m8